OpenStudy (anonymous):

How to make this integral ∫dP/(P- P^2 ) thanks in advance!!!

OpenStudy (anonymous):

\[\int\frac1{p(1-p)}dp = \int\frac{1}{p}dp + \int\frac{1}{1-p}dp = \ln p + \ln (1-p) +c\]

OpenStudy (dumbcow):

except the last term should be -ln(1-p) u = 1-p du = -dp ...

OpenStudy (anonymous):

Oh yeah, thanks chris.

OpenStudy (anonymous):

thank you so much!